Find the equation of the line passing through (-1,2) and perpindicular to y=3x+5
Answer by stanbon(75887)

Find the equation of the line passing through (-1,2) and perpindicular to y=3x+5
-------
The given line has a slope of 3.
-----
All perpendicular lines must have a slope of -1/3
-----------
Find equation of line with slope = -1/3 passing thru (-1,2):
y = mx + b
2 = (-1/3)(-1) + b
2 = (1/3) + b
b = 5/3
--------------
Equatiion:
y = (-1/3)x + (5/3)
========================
Cheers,
Stan H.
===========